OpenStudy (anonymous):

Find the inverse formula of: y = f(t) = (2t)^(1/2) + 7

OpenStudy (anonymous):

Here is what I have so far: t = square root of (2y) + 7. t - 7 = square root of 2y (t-7)(t-7) = 2y t^(2) - 14t + 49 = 2y (t^(2)-14t + 49)/2 = y The website is saying the variable 't' is not defined in this context. I tried X, but it didn't work either

OpenStudy (anonymous):

y=(2t)^1/2+7 Inter changing x and t t=(2y)^1/2+7 t-7=(2y)^1/2 t^2-14t+49=2y y=1/2t^2-7t+49/2

OpenStudy (anonymous):

lol.... got the same

OpenStudy (anonymous):

why do you have 1/2t^(2)?